# CALF: Aligning LLMs for Time Series Forecasting via Cross-modal Fine-Tuning

### v2

@@ 4.3 Few/zero-shot Learning
We conduct few-shot experiments on four ETT datasets. Specifically, for each dataset, we utilize only the first 10% of the training data. This constrained data scenario presents a considerable challenge, testing the ability of the model to learn effectively with limited information. Tab. 3 demonstrates that our method outperforms other baselines, highlighting its robustness in the few-shot setting. Compared with GPT4TS and PatchTST, our method achieves an average reduction of 8% and 9%, respectively.
Models CALF TimeLLM GPT4TS PatchTST Crossformer FEDformer TimesNet MICN DLinear TiDE
[12]
[13]
[6]
[5]
[9]
[4]
[17]
[7]
[8]
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1
0.504 0.462 0.636 0.512 0.608 0.500 0.557 0.483 1.340 0.848 0.696 0.572 0.673 0.534 0.970 0.674 0.567 0.499 0.515 0.469
ETTm2
0.302 0.330 0.348 0.343 0.303 0.336 0.295 0.334 1.985 1.048 0.356 0.392 0.321 0.354 1.073 0.716 0.329 0.382 0.303 0.337
ETTh1
0.644 0.541 0.765 0.584 0.689 0.555 0.683 0.546 1.744 0.914 0.750 0.607 0.865 0.625 1.405 0.814 0.647 0.552 0.779 0.604
ETTh2
0.419 0.427 0.589 0.498 0.579 0.497 0.550 0.487 3.139 1.378 0.553 0.525 0.476 0.463 2.533 1.158 0.441 0.458 0.421 0.428
Table 3: Few-shot learning results on 10% training data of ETT datasets. All the results are averaged from 4 different prediction lengths $H\in\{96,192,336,720\}$. Sec. F.3 shows the full results.
@@
Going beyond few-shot scenarios, we further delve into zero-shot learning, where LLMs demonstrate their prowess as adept and intuitive reasoners. In this setting, models trained on one dataset $\blacklozenge$ are evaluated on an entirely different dataset $\bigstar$, without any further training. As shown in Tab. 4, our method stands out for its exceptional performance, surpassing GPT4TS and PatchTST by 4% and 9% respectively. This indicates that our approach significantly enhances the model’s capability for effective learning transfer across different domains.
Models CALF TimeLLM GPT4TS PatchTST Crossformer FEDformer TimesNet MICN DLinear TiDE
[12]
[13]
[6]
[5]
[9]
[4]
[17]
[7]
[8]
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
h1 $\rightarrow$ m1
0.755 0.574 0.847 0.565 0.798 0.574 0.894 0.610 0.999 0.736 0.765 0.588 0.794 0.575 1.439 0.780 0.760 0.577 0.774 0.574
h1 $\rightarrow$ m2
0.316 0.355 0.315 0.357 0.317 0.359 0.318 0.362 1.120 0.789 0.357 0.403 0.339 0.370 2.428 1.236 0.399 0.439 0.314 0.355
h2 $\rightarrow$ m1
0.836 0.586 0.868 0.595 0.920 0.610 0.871 0.596 1.195 0.711 0.741 0.588 1.286 0.705 0.764 0.601 0.778 0.594 0.841 0.590
h2 $\rightarrow$ m2
0.319 0.360 0.322 0.363 0.331 0.371 0.420 0.433 2.043 1.124 0.365 0.405 0.361 0.390 0.527 0.519 0.496 0.496 0.321 0.364
[…]
Time (s) MSE / MAE
ETTm1 ETTh1 ECL Traffic Weather ETTm1 ETTh1 ECL Traffic Weather
GPT4TS [13] 626 81 8274 15067 596 0.329 / 0.364 0.376 / 0.397 0.185 / 0.272 0.468 / 0.307 0.182 / 0.223
Time-LLM [12] 1476 314 33209 62412 1262 0.359 / 0.381 0.398 / 0.410 0.204 / 0.293 0.536 / 0.359 0.195 / 0.233
CALF (Ours) 135 27 251 614 123 0.323 / 0.349 0.369 / 0.389 0.145 / 0.238 0.407 / 0.268 0.164 / 0.204
Table 5: Comparison of different LLM-based time series forecasting methods in terms of computation time and performance (MSE/MAE) across various datasets. The input and predict length are both set to 96.
4.4 Efficiency Analysis
We conduct experiments on five datasets: ETTm1, ETTh1, ECL, Traffic, and Weather. The input and prediction lengths are both set to 96. As shown in Tab. 5, our proposed CALF shows significant improvements in both efficiency and accuracy compared with other LLM-based methods. We also provide theoretical complexity analysis for various Transformer-based methods in Appendix E.
5 Abaltion Study
@@
The feature regularization loss $\mathcal{L}_{feature}$ aligns the intermediate features between the textual source branch and the temporal target branch, while the output consistency loss $\mathcal{L}_{output}$ ensures output coherence across modalities. The supervised loss $\mathcal{L}_{sup}$ directly guides learning with ground truth data. We analyze the specific effects of each proposed loss function as detailed in Sec. 5. Employing only the supervised loss resulted in MSE/MAE of 0.446/0.438 on ETTh1 and 0.263/0.286 on Weather, respectively. The addition of feature regularization loss $\mathcal{L}_{feature}$ or output consistency loss $\mathcal{L}_{output}$ led to incremental improvements, with the best performance observed when all three losses were combined, achieving the lowest MSE and MAE on both datasets.
@@
Ablation on different loss functions on ETTh1 and Weather datasets.
@@
6 Discussion
Difference form Other Work.
One concurrent work [12] also considers cross attention to extracting knowledge from the word embedding layer, and we would like to clarify the difference to emphasize our contribution. First, the existing method uses cross-attention to generate embeddings and combines them with prompt prefixes as input to frozen LLMs, while our CALF aims to generate aligned textual tokens as the input of the textual modal branch for subsequent cross-modal distillation. Second, previous work introduces linear weight $W\in\mathbb{R}^{|\mathcal{A}|\times d}$ to learn text prototype during training. However, given the huge word space $|\mathcal{A}|$, this solution can lead to significant costs, while our approach uses an offline manner to generate synonym clusters, which guarantees efficiency.
@@ 6 Discussion
Limitations and Future Works.
Our input alignment method relies on implicit alignment, which may not fully leverage the explicit textual reasoning capabilities inherent in LLMs [44]. Existing methods use explicit text merely as prior knowledge [12], missing opportunities for deeper integration. Future works should focus on seamlessly incorporating explicit textual information into time series analysis through improved pre-training techniques or advanced representation methods.
7 Conclusion

### v3

@@ Results.
We conduct few-shot experiments on four ETT datasets. Specifically, for each dataset, we utilize only the first 10% of the training data. This constrained data scenario presents a considerable challenge, testing the ability of the model to learn effectively with limited information. Tab. 3 demonstrates that our method outperforms other baselines, highlighting its robustness in the few-shot setting. Compared with TimeLLM and PatchTST, our method achieves an average reduction of 8% and 9%, respectively.
Models h1 $\rightarrow$ m1 h1 $\rightarrow$ m2 h2 $\rightarrow$ m1 h2 $\rightarrow$ m2
MSE MAE MSE MAE MSE MAE MSE MAE
TiDE 0.774 0.574 0.314 0.355 0.841 0.590 0.321 0.364
DLinear 0.760 0.577 0.399 0.439 0.778 0.594 0.496 0.496
MICN 1.439 0.780 2.428 1.236 0.764 0.601 0.527 0.519
TimesNet 0.794 0.575 0.339 0.370 1.286 0.705 0.361 0.390
FEDformer 0.765 0.588 0.357 0.403 0.741 0.588 0.365 0.405
Crossformer 0.999 0.736 1.120 0.789 1.195 0.711 2.043 1.124
PatchTST 0.894 0.610 0.318 0.362 0.871 0.596 0.420 0.433
GPT4TS 0.798 0.574 0.317 0.359 0.920 0.610 0.331 0.371
TimeLLM 0.847 0.565 0.315 0.357 0.868 0.595 0.322 0.363
CALF 0.755 0.574 0.316 0.355 0.836 0.586 0.319 0.360
[…]
@@
Going beyond few-shot scenarios, we further delve into zero-shot learning, where LLMs demonstrate their prowess as adept and intuitive reasoners. In this setting, models trained on one dataset $\blacklozenge$ are evaluated on an entirely different dataset $\bigstar$, without any further training. As shown in Tab. 4, our method stands out for its exceptional performance, surpassing TimeLLM and PatchTST by 4% and 9% respectively. This indicates that our approach significantly enhances the model’s capability for effective learning transfer across different domains.
Dataset GPT4TS Time-LLM CALF
Time (s) MSE / MAE Time (s) MSE / MAE Time (s) MSE / MAE
ETTm1 626 0.329 / 0.364 1476 0.359 / 0.381 135 0.323 / 0.349
ECL 8274 0.185 / 0.272 33209 0.204 / 0.293 251 0.145 / 0.238
Traffic 15067 0.468 / 0.307 62412 0.536 / 0.359 614 0.407 / 0.268
Weather 596 0.182 / 0.223 1262 0.195 / 0.233 123 0.164 / 0.204
Table 5: Comparison of different LLM-based time series forecasting methods in terms of computation time and performance (MSE/MAE) across various datasets.
Efficiency Analysis
We conduct experiments on four datasets: ETTm1, ECL, Traffic, and Weather. The input and prediction lengths are both set to 96. As shown in Tab. 5, our proposed CALF shows significant improvements in both efficiency and accuracy compared with other LLM-based methods, largely due to treating each channel sequence as a token, employing an efficient fine-tuning strategy, and requiring only a single time branch during inference.
Ablation Study
@@
The feature regularization loss $\mathcal{L}_{feature}$ aligns the intermediate features between the textual source branch and the temporal target branch, while the output consistency loss $\mathcal{L}_{output}$ ensures output coherence across modalities. The supervised loss $\mathcal{L}_{sup}$ directly guides learning with ground truth data. We analyze the specific effects of each proposed loss function as detailed in Tab. 6. Employing only the supervised loss resulted in MSE/MAE of 0.446/0.438 on ETTh1 and 0.263/0.286 on Weather, respectively. The addition of feature regularization loss $\mathcal{L}_{feature}$ or output consistency loss $\mathcal{L}_{output}$ led to incremental improvements, with the best performance observed when all three losses were combined, achieving the lowest MSE and MAE on both datasets.
@@
Table 6: Ablation on different loss functions on ETTh1 and Weather datasets.
Discussion
Difference from Other Work.
Existing LLM-based methods, such as TimeLLM (Jin et al. 2024) and TEST (Sun et al. 2024), also consider the alignment between the time-modal input and the text-modal parameters by employing the cross-attention layer or contrastive learning at the input side. However, we point out that this simple scheme is not sufficiently powerful thus leading to partial alignment (Fig. 1) as well as weak generalization performance (Tab. 3, Tab. 4). By contrast, we identify the alignment problem and further propose a multi-level cross-modal fine-tuning framework to achieve finer-grained alignment. Moreover, although the previous approaches adopt one cross-attention layer as an intuitive solution, it can raise significant computational costs due to the huge alphabet size of LLM. In this work, we introduce the offline PCA to generate synonym clusters from the huge alphabet as an alternative, allowing high efficiency while achieving better performance.
@@ Discussion
Conclusion